OpenStudy (anonymous):

If the dimensions of a rectangle become one-fourth of the original dimensions, then area of the new rectangle will be _______of the area of the original rectangle.

OpenStudy (anonymous):

one sixteenth

OpenStudy (saifoo.khan):

The formula is something like, \[A = L^2\]

OpenStudy (saifoo.khan):

So, \[A = (\frac{1}{4})^2\]

OpenStudy (anonymous):

A=L*B A'=(L/4)*(B/4)=(LB)/16=A/16

jimthompson5910 (jim_thompson5910):

A = LW A' = (L/4)*(W/4) A' = (LW)/16 So the area of the new rectangle A' is 1/16 of the area of the original rectangle A

OpenStudy (anonymous):

Thank you all!!